OpenStudy (anonymous):

simplify the expression: sinx+cosx/sinxcosx I know to separate them first into sinx/sinxcosx + cosx/sinxcosx , but then i get lost after that

OpenStudy (anonymous):

Sin[x]/(sin[x]cos[x])=1/Cos[x]=Sec[x]....Cos[x]/(Cos[x]Sin[x])=1/Sin[x]=Csc[x]... So you would have Sec[x]+Csc[x]

OpenStudy (anonymous):

where u get 1/cosx from when u solved for secx?

OpenStudy (anonymous):

Sec[x] is 1/cosx

OpenStudy (anonymous):

or cosx=1/secx

OpenStudy (anonymous):

how did u get sinxcosx to equal 1/cosx?

OpenStudy (anonymous):

Reciprocal Identities are the following: Sec[x]=1/Cos[x] and Csc[x]=1/sin[x] and Cot[x]=1/Tan[x]..

OpenStudy (anonymous):

well the problem asks to simplify (sin[x]+cos[x])/(sin[x]cos[x]), correct?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Okay, you made the correct step by getting sin[x]/sin[x]cos[x]+cos[x]/sin[x]cos[x]... From here the sin[x] cancels out on the left and the cos[x] cancels on the right to give 1/cos[x]+1/sin[x]

OpenStudy (anonymous):

oh! i see now ^^ thx!

OpenStudy (anonymous):

You welcome.